OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

Mehek (mehek14):

combine like terms -3x + 6x =?

Mehek (mehek14):

\(-3x - 8 + 6x = 7\) combine like terms \(3x - 8 = 7\) add 6 \(3x = 15\) divide by 3 \(x = 15\)

OpenStudy (anonymous):

thanks @Mehek14 just saved my life